OpenStudy (anonymous):

How do you change powers with negative exponents to powers with positive exponents?

OpenStudy (anonymous):

\[\frac{x^{-5}}{1} = \frac{1}{x^5}\]

OpenStudy (anonymous):

in words please. i know what to do but dont know who to word it???

OpenStudy (anonymous):

The proof is \[ \frac{x^{-5}}{1}\cdot \frac{x^5}{x^5 }=\frac{x^{-5}x^5}{x^5}=\frac{x^0}{x^5}=\frac{1}{x^5}\]

OpenStudy (anonymous):

Ok, you could say you multiply the term by its inverse and then divide by the inverse to keep the expression the same....

OpenStudy (anonymous):

You end up with 1 divided by the original element's inverse which is what you are doing...